OpenStudy (anonymous):

Determine two pairs of polar coordinates for the point (4, -4) with 0° ≤ θ < 360°. Can somebody please help? I don't understand.

OpenStudy (anonymous):

@emcrazy14 @deinomios @norasparkle4

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@experimentX

OpenStudy (anonymous):

@DanielM_113

OpenStudy (anonymous):

Do you know the equations for polar coordinates?

OpenStudy (anonymous):

I don't think so

OpenStudy (anonymous):

Do you know what they mean?

OpenStudy (anonymous):

Ok so now what do i do?

OpenStudy (anonymous):

Now you convert the given point to polar coordinates if you understand what polar coordinates mean. Point(4,-4) or x=4 and y=-4, turn that into r and \(\theta\).

OpenStudy (anonymous):

so is the answer (4 sqrt 2, 135 degrees), (-4 sqrt 2, 315 degrees)?

OpenStudy (anonymous):

I am not sure. Let me check.

OpenStudy (anonymous):

ok thanks so much! I really appreciate it!

OpenStudy (anonymous):

Yes, your answer is correct.

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

You are welcome.

OpenStudy (anonymous):

Could you help me with one more?

OpenStudy (anonymous):

@DanielM_113

OpenStudy (anonymous):

Yes, what is the question?

OpenStudy (anonymous):

Find all polar coordinates of point P where P = (9, pi/5) @DanielM_113

OpenStudy (anonymous):

You use the same equations as before, but now x=9 and y=\(\pi\)/5

OpenStudy (anonymous):

Is the answer (9, pi/5 + 2nπ) or (-9, pi/5 + 2nπ)?

OpenStudy (anonymous):

\[ r=\sqrt{9^2+\frac{\pi}{5}}\] hmm, if that is supposed to be the answer I misunderstood the question. If the point P is already in polar coordinates than your answer is correct.

OpenStudy (anonymous):

I think it is. Thank you for your help!!

OpenStudy (anonymous):

The way you understood makes much more sense, I must have understood the problem wrong.

OpenStudy (anonymous):

No problem.